OpenStudy (anonymous):

Find x or y (3,y) (1,3y)

OpenStudy (anonymous):

Do you know your rate of change?

OpenStudy (anonymous):

Nope I do not know how to figure that out

OpenStudy (anonymous):

wait there isn't enough numbers to do that anyways I think

OpenStudy (anonymous):

Correct. So, X is already given, so we must find Y. To do this, We must identify the rate of change. Is there a graph that goes with this question?

OpenStudy (anonymous):

noope just those 2 points

OpenStudy (anonymous):

Okay, then we will use Point-Slope form. Do you know this?

OpenStudy (anonymous):

no i don't

OpenStudy (anonymous):

(Y-Y1)=M(X-X1)

OpenStudy (anonymous):

oh yes i do know this

OpenStudy (anonymous):

(y-y) = 0(x-3)

OpenStudy (anonymous):

how do we solve that

OpenStudy (anonymous):

@highschoolmom2010 @HelpBlahBlahBlah

OpenStudy (anonymous):

@triciaal